OpenStudy (anonymous):

At the farmer's market, there is a scale that measures the weight of fruit and vegetables. The needle is 16 cm long and it rotated 68° when Danielle placed six oranges on it. How far did the needle travel as it rotated?

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

a: (272)(pi) / 45 b: (45)(pi) / 272 c: 1088(pi) d: 544(pi)

OpenStudy (anonymous):

@Jhannybean

OpenStudy (jhannybean):

Let's say arc length = L. Then \[L=r\theta\]You need to change the angle of rotation to radians, and to do that, you solve for: \[\theta = 60^\circ\cdot \frac{\pi}{180^\circ}\]Plug in all your values and solve for \(L\)

OpenStudy (anonymous):

im confused

OpenStudy (jhannybean):

What part are you confused with?

OpenStudy (anonymous):

i dont understand how to change into radians, and/or eventually get the answer

OpenStudy (jhannybean):

 You've learned that arc length = radius * angle measurement right? That is equivalent of saying \[s=r\theta\]Instead of \(s\) I'm using \(L\) for length.

OpenStudy (jhannybean):

The formula to convert from radian to degrees and degrees to radians is: \(\dfrac{ 180^\circ}{\pi}\) To change a degree to a radian, or vise versa, all you have to do is know what you want to convert to. If you want to convert to radians, \(\pi\) will be on top. If you're given a radian measurement and want to convert to a degree measurement, put \(180^\circ\) on top.

OpenStudy (anonymous):

oohk okaay, then how do you finish it?

OpenStudy (jhannybean):

In this case, we want to convert to radians. Therefore we keep \(\pi\) on top. \[68^\circ \cdot \frac{\pi}{180^\circ}\]This cancels out the degrees and leaves us in radians.

OpenStudy (jhannybean):

Calculate \[L=16~cm \cdot \left(\frac{60^\circ}{180^\circ} \cdot \pi\right)\] and you'll have your answer.

OpenStudy (jhannybean):

I will not solve it for you.

OpenStudy (anonymous):

i got 16(pi)/3

OpenStudy (anonymous):

nvm i got A: 272(pi) / 45